A company car that has a seating capacity of seven is to be used by seven employees who have formed a car pool. If only three of these employees can drive, how many possible seating arrangements are there for the group?
Answer by ikleyn(52780)

.

Any of the three who can drive, can seat on the driver's seat,  giving 3 possible options.


For any such driver, the rest can be arranged in 6! = 720 ways.


So, the total number of arrangements is 3*6! = 3*720 = 2160.    ANSWER

Solved.